OpenStudy (anonymous):

write a quadratic equation in the variable x having the given numbers as solutions. type the equation in standard form ax^2+bx+c=0 the solutions are -11,5

OpenStudy (saifoo.khan):

(x+11)(x-5) = 0 FOIL

OpenStudy (anonymous):

shouldn't it be (x-11)(x+5)=0?

OpenStudy (saifoo.khan):

No, b/c the solutions are x = -11 x + 11 = 0

OpenStudy (anonymous):

ok so when I foil it then what?

OpenStudy (anonymous):

is it x^2+7x-55?